### v2

# ShowHowTo: Generating Scene-Conditioned Step-by-Step Visual Instructions

Abstract
The goal of this work is to generate step-by-step visual instructions in the form of a sequence of images, given an input image that provides the scene context and the sequence of textual instructions. This is a challenging problem as it requires generating multi-step image sequences to achieve a complex goal while being grounded in a specific environment. Part of the challenge stems from the lack of large-scale training data for this problem. The contribution of this work is thus three-fold. First, we introduce an automatic approach for collecting large step-by-step visual instruction training data from instructional videos. We apply this approach to one million videos and create a large-scale, high-quality dataset of 0.6M sequences of image-text pairs. Second, we develop and train ShowHowTo, a video diffusion model capable of generating step-by-step visual instructions consistent with the provided input image. Third, we evaluate the generated image sequences across three dimensions of accuracy (step, scene, and task) and show our model achieves state-of-the-art results on all of them. Our code, dataset, and trained models are publicly available.
Figure 1: Given an input image (left) and ordered step-by-step textual instructions for a task (top), ShowHowTo generates an image sequence of visual instructions. Rows 1 and 2 demonstrate the generation of visual instructions for two recipes starting from the same input image. Rows 2 and 3 show the generation of visual instructions for the same recipe but conditioned on different input images. ShowHowTo generates scene-consistent (e.g., consistency in the person and cutting board) and temporally consistent image sequences (e.g., the bowl of tortilla chips or plate of chicken skewers) that faithfully capture the instructions (e.g., cutting, frying, brushing, adding etc.).
11footnotetext: Czech Institute of Informatics, Robotics and Cybernetics at the Czech Technical University in Prague.22footnotetext: School of Computer Science, Machine Learning and Computer Vision (MaVi) Research Group, University of Bristol, UK.33footnotetext: Mohamed bin Zayed University of Artificial Intelligence.
1 Introduction
With the immense success of large vision-language models and the rise of wearable devices, we rapidly approach the era of personalized visual assistants. This technology promises to help us in a variety of everyday tasks and numerous scenarios, such as preparing a Michelin-star dish, taking care of plants, or fixing a bicycle. Unlike generic instructional videos, visual assistants will provide guidance and feedback for specific environments and task variations.
Besides being useful for people, automated visual guidance has been recently explored and shown to be beneficial in robotics. For example, [8, 32, 43, 65] generate images of intermediate goals and use them as guidance for manipulation policies. Other recent methods [7, 18, 38, 57] derive robotics policies from videos specifically generated for target tasks and environments.
When comparing the ability to generate textual step-by-step instructions vs. generating visual instructions, one can see a sharp contrast. State-of-the-art LLMs can reliably provide personalized step-by-step text-only instructions. However, translating such instructions into images and videos still presents considerable challenges. This is because current video generation models, despite their impressive progress over the past years, focus on producing relatively short clips [13, 46, 49, 63] whereas image generation models only produce one image at a time.
Recent attempts to generate visual instructions either synthesize a single step [35, 36, 53] or are not contextualized to the user’s specific environment [41, 45, 11]. In other words, such methods may generate plausible images for each step, however, such images will represent arbitrary settings and can feature tools or ingredients unavailable to the user. In robotics, temporally inconsistent guidance may imply physically implausible demonstrations, resulting in unsuccessful learning of policies. To address this issue, we focus on generating step-by-step visual instructions conditioned on an input image from the user, which we assume showcases their starting position—the environment, ingredients, tools, etc., as illustrated in Figure 1.
This paper makes the following contributions. (1) We introduce the problem of generating a sequence of visual instructions conditioned on an input image. (2) We introduce a fully automatic approach to collect step-by-step visual instruction training data from in-the-wild instructional videos, creating a large-scale, high-quality dataset of 0.6M step-by-step instruction sequences of 4.5M image-text pairs. (3) We train a video diffusion model capable of generating sparse step-by-step visual instructions consistent with the input image. (4) We evaluate our generated sequences across three aspects (step, scene, and task) and show our model achieves state-of-the-art results on all of them.
2 Related Work
Figure 2: Our automatic approach for creating the ShowHowTo dataset—a large-scale instructional dataset consisting of step-by-step instruction sequences of image-text pairs to perform diverse HowTo tasks. Examples of step-by-step textual instructions and the corresponding frames are highlighted in green.
Datasets of visual instructions. The scale and quality of the training data play a key role in visual instruction generation. Many available datasets combine instructional or egocentric videos and manual temporal annotation of individual steps [51, 54, 70, 2] or use professional illustrations [62]. Yet the requirement of manual annotations makes these sources of data hard to scale to novel tasks and environments. To alleviate the need for manual annotations, self-supervised methods have been developed to automatically obtain key steps from in-the-wild videos [59, 39, 58, 20, 52]. These key steps can then be used for visual instruction generation [53]. Recently, the improved capabilities of large language models [19, 1, 17, 33, 3] allowed for solely using video narrations to produce temporal captions, key steps, and instructions [37, 50, 36]. We build on these works to automatically obtain key steps from videos; however, in contrast to the related works, our dataset is constructed completely automatically, is composed of individual instruction frames instead of temporal intervals, and contains significantly fewer errors. Additionally, we focus on the entire domain of instructional videos, not only cooking.
Conditional video generation. Recently, diffusion models [27] have seen a surge in popularity for generative tasks, including video generation [64, 40]. Initial works extended a U-Net model using space-time factorization for the generation of videos in pixel space [29, 28, 30]. With the large cost of generating video, others [14, 69, 26, 10, 22, 64, 40] instead use an auto-encoder to model videos within the latent space, reducing significantly the number of parameters and memory requirements. Video generation models are often conditioned on textual prompts [23, 15, 29, 28, 67, 22], images that act as initial frames [60, 68], or both to generate a sequence of frames [9, 6, 24, 34, 30, 69, 10, 16, 56]. It has also been shown that for temporal consistency, a combination of both image conditioning and textual conditioning is important [55, 56]. However, these methods focus on relatively short video clips and are not able to generate long multi-step sequences of fine-grained instructions that take minutes to execute.
Generating step-by-step instructions. Both step-by-step visual instructions and continuous videos consist of sequences of frames, yet the instruction sequences differ from the videos significantly. While videos often contain only small pixel-level frame-to-frame changes [61], visual instruction sequences often contain large semantic (e.g., raw $\rightarrow$ cooked) and viewpoint (e.g., inside $\rightarrow$ outside) changes from one key frame to another [41]. Obtaining sufficient training data for visual instructions presents a significant challenge. Therefore, Phung et al. [45] generate step-by-step visual instructions using a pretrained text-conditioned image diffusion model with shared attention across steps to ensure consistency in the generated image sequences, while Menon et al. [41] use illustrations drawn by artists from WikiHow [62] as the training data for text-to-image-sequence generation. Other works using image-conditioned models can generate step-by-step sequences by iterative generation [11, 53, 36, 35]. In contrast, our method generates step-by-step visual instructions all at once, attending across steps to generate the full image sequence, including the input, which results in superior quality and consistency.
3 Building Large-Scale ShowHowTo Dataset
Learning to generate visual instructions requires a large-scale dataset that captures the rich diversity of real-world tasks and their step-by-step execution. However, manually creating such a dataset is prohibitively expensive and time-consuming, limiting the dataset’s scale and coverage. We address this challenge by introducing an automated approach that leverages the natural alignment between narrations and visual demonstrations in instructional videos from the web to mine high-quality sequences of image-text pairs.
Using our proposed approach, we construct a large-scale dataset containing over half a million instruction sequences of image-text pairs spanning 25,026 diverse HowTo tasks. These sequences cover diverse domains including cooking (e.g., make strawberry crumb bars, prepare an avocado margarita), home improvement (e.g., stain a cabinet, create a tire garden), assembly (e.g., set up a 10$\times$10 tent, tie a ring sling), DIY crafts (e.g., make a bracelet, make a fairy glow jar) and many more. We note that our data collection approach does not require any manual annotation, which is an important aspect to enable scaling.
3.1 Automatic Dataset Collection
Our approach takes as input a narrated instructional video for a specific task. First, it extracts a sequence of key steps in the form of concise, free-form textual instructions from the video’s narration. Then, it associates each step with the corresponding keyframe in the video. The output is an ordered sequence of image-text pairs. This is a very challenging task due to the high level of noise, the possible misalignment of the narration and the visual content as well as the sheer variety of visual appearance and the spoken natural language in the input internet videos.
To tackle these challenges, we design a four-stage approach, illustrated in Figure 2: (1) The narration of the input internet instructional video is transcribed into sentences with corresponding timestamps. (2) The transcribed narration is verified to be instructional and removed if not. (3) The key instruction steps are extracted from the transcript along with their approximate temporal bounds. (4) A representative frame for each instruction step is selected through cross-modal alignment. By applying this approach to videos from HowTo100M [42], we obtain 578K high-quality sequences of image-text pairs with approximately eight steps per video on average.
Formally, we define our dataset as a collection of instruction sequences of image-text pairs. Each sequence $\{(I_{i},\tau_{i})\}_{i=0}^{n}$ represents an ordered set of steps required to accomplish a specific task $\mathcal{T}$. It consists of pairs of images $I_{i}$ and the corresponding natural language instructions $\tau_{i}$, with $n$ denoting the number of steps in the sequence. Next, we describe the four stages in detail.
Speech transcription. Accurate transcription (ASR) of spoken narrations is a key strength of our approach, as these transcripts capture the instructor’s step-by-step guidance that we use to align with the video. HowTo100M provides 1.2 million web instructional videos, but we forego the original transcripts generated using the YouTube API due to noise [25, 37]. Instead, we use WhisperX [5], a state-of-the-art speech recognition model, to obtain high-quality transcriptions with accurate timestamps from videos. We provide comparisons of the original transcripts and the improved ones in Figure 18.
Filtering of irrelevant videos. We find that many HowTo100M videos are non-instructional, containing product reviews, vlogs, movie clips, etc. This noise may stem from the original data collection process, which relied on keyword-based web crawling and is susceptible to false positives due to inaccurate metadata. We leverage video transcripts as a strong signal for identifying instructional content and use a recent LLM (Llama 3.1 [19]) to filter the videos. We verify the reliability of this process through the evaluation on a labeled subset. Detailed analysis, qualitative results, and the prompts used for querying the LLM are provided in Appendix A and Figure 19.
Step extraction. We observe that, in instructional videos, the key steps necessary to achieve a particular task are very often mentioned in the narration, even if they are not well-aligned with what is shown in the frame [25]. Building on this, we prompt an LLM to extract the instructional steps from the narration transcripts in the format of WikiHow step-by-step guides, providing exemplars in the prompt to guide the extraction. Somewhat surprisingly, the LLM not only correctly extracts the key steps from the transcripts, but the model is also able to associate each step with the correct temporal intervals from the transcript, even if the step spans over multiple narrations. See Figure 2 for an example, and Appendix A and Figure 20 for additional details and the prompt used.
Cross-modal frame alignment. For each instructional step, our goal is to identify a single representative frame that best demonstrates the instruction visually. While contrastive models [47, 66, 21] can align text instructions with frames, we observe that naive text matching across thousands of video frames leads to noisy results. Therefore, we limit the alignment to the identified instruction step temporal interval, expanded by $\epsilon=15$ seconds to allow for some level of misalignment between the narrations and visual demonstrations [25]. Given these expanded intervals, we compute text-frame similarity scores using DFN-CLIP [21] and select the best alignment that satisfies the temporal ordering of the steps. We provide more details about the matching process in Appendix A.
3.2 Dataset Statistics
In total, after filtering, the dataset contains 578K unique sequences of image-text pairs, with a total of 4.5M steps, averaging 7.7 ($\pm$ 2.8) steps per sequence, and 11.4 ($\pm$ 4.7) words per step. The sequence lengths vary from 1 to 26 steps, with $97.6\%$ of sequences being 2 to 16 steps long. From task information provided by HowTo100M, the dataset contains instructions for 25K HowTo tasks across several categories, such as cooking, home and garden improvement, vehicles, personal care, health, and more. We include a comparative table to other datasets and further dataset analysis in the appendix.
4 ShowHowTo Model and Training Procedure
Figure 3: Model architecture. Given an input frame $I_{0}$ (left) and a variable number of text instructions $\tau_{i}$ describing each step, our diffusion model generates visual instructions $\hat{I}_{i}$ that correctly follow the prompts $\tau_{i}$ and are consistent with the input image $I_{0}$.
Given a user-provided image $I_{0}$, such as a photo of ingredients or tools on a table, our goal is to generate a sequence of images $\{\hat{I}_{i}\}_{i=1}^{n}$ of any length $n$ based on the number of required steps, that guides the user to achieve an intended task $\mathcal{T}$, such as a cooked chicken tikka masala dish. Our goal is to generate images $\hat{I}_{i}$ to match the user-provided context, i.e., to be grounded in the user’s environment by utilizing the specific objects, tools, and workspace from the input image $I_{0}$. We achieve this goal by training a diffusion model conditioned on the input image $I_{0}$ along with the step-by-step textual instructions $\{\tau_{i}\}_{i=0}^{n}$ that fulfill the intended task $\mathcal{T}$ in any number of steps $1\leq n\leq 15$.
We build on recent progress in diffusion models for video generation [56]. However, there are the technical challenges of (a) how to inject the multi-step instruction guidance and (b) how to generate variable length sequences. We address these challenges in the next paragraphs.
Architecture. Our model, shown in Figure 3, is based on a latent video diffusion model [56] composed of a U-Net encoder and decoder, each with interleaving spatial and temporal attention layers. The input image $I_{0}$ is projected into the latent space via the VAE encoder $\mathcal{E}$, and is concatenated to each frame of the random noise $z_{i}$ to form the model’s input. The U-Net progressively denoises the input latent sequence while attending to all images in the sequence to ensure the generated images are temporally consistent and aligned with the input image. For better conditioning on the input image, the U-Net also contains cross-attention layers that attend to a feature representation of the input image directly. To guide the generation process to the desired visual instruction sequence, each frame $i$ in the sequence attends to its prompt $\tau_{i}$ via cross-attention layers of the U-Net. In the ablations, we show that separate text conditioning for each frame in the sequence is instrumental for generating high-quality step-by-step visual instructions.
Training. We initialize the model from the pretrained checkpoint [56] trained on WebVid10M [4] for image animation and fine-tune the entire U-Net weights on our dataset. In contrast to training on videos, where the output video is commonly of a fixed length (e.g., 16 frames in the case of [56]), step-by-step instructions have a variable sequence length. To ensure our model can generate variable-length sequences, we vary the sequence length during training by sampling from our sequences. For efficient computation, the length is varied over different batches but is kept the same across all samples in a single batch. If the dataset sequence is longer than the desired length $k$, we randomly sample the starting frame and use the next consecutive $k$ frames as the model’s target. We verify and further discuss these choices in Section 5.3 and provide the implementation details in Appendix C.
5 Experiments
We first introduce our evaluation setup in Section 5.1. In Section 5.2, we compare our method to current state-of-the-art quantitatively on the test set and through a user study. Section 5.3 analyzes key designs of our method through ablation studies. Finally, Section 5.4, showcases qualitative results of our method. For implementation details and additional results, see the appendix.
5.1 Evaluation details
Dataset. We construct train and test splits from our dataset of 578K samples. Our test set comprises 3,964 sequences from 200 tasks covering the distribution over task categories in the full dataset. To ensure sample quality, we prioritize samples with high DFN-CLIP alignment scores as measured in our dataset creation pipeline (Section 3). For zero-shot evaluation of our method, we also use a random subset of 1442 non-illustrated instructional sequences [^1] See the project website for the list of selected sequences. from the WikiHow-VGSI dataset [62] as an additional test set.
Evaluation metrics. We evaluate our model by measuring the correctness and consistency of the generated visual instructions using similar metrics as Menon et al. [41]. We describe the used metrics next with full details available in Appendix D. (1) Step Faithfulness [41] measures whether each generated image $\hat{I}_{i}$ correctly depicts its corresponding text instruction $\tau_{i}$. It is computed as the zero-shot accuracy of the DFN-CLIP model where the generated image $\hat{I}_{i}$ is classified into classes $\{\tau_{i}\}_{i=0}^{n}$ of all text instructions of the sequence. (2) Scene Consistency measures whether the generated image $\hat{I}_{i}$ consistently captures the scene from the input image $I_{0}$ (e.g., the same utensils are used on the same kitchen countertop as in the input image). Intuitively, a generated image $\hat{I}_{i}$ is considered scene-consistent if it visually matches any frame from its source video $\{I_{i}\}_{i=1}^{n}$ (excluding the input image to avoid trivial copy solution). Therefore, for each generated image $\hat{I}_{i}$, the most similar image according to the DINOv2 [44] score is retrieved from the test set images. The metric then measures if the retrieved image is from the same video as the input. (3) Task Faithfulness measures how well the generated sequence $\{\hat{I}_{i}\}_{i=1}^{n}$ represents its intended task. It is measured as the zero-shot accuracy of the DFN-CLIP model where the generated sequence’s averaged feature vector is classified into all 200 test set tasks. In contrast to Menon et al. [41], the generated sequences are evaluated holistically rather than per-step, as often steps are not unique to a task (e.g., “knead dough” step appears in both “Make sourdough bread” and “Make pizza” tasks), and the classification is done into all test set tasks rather than a small random subset, providing a more robust evaluation.
5.2 Comparison with the State-of-the-Art
ShowHowTo WikiHow [62]
Method Step Scene Task Step Scene
Faithf. Consist. Faithf. Faithf. Consist.
(a) InstructPix2Pix [12] 0.25 0.17 0.25 0.32 0.12
(b) AURORA [35] 0.25 0.33 0.24 0.33 0.15
(c) GenHowTo [53] 0.49 0.13 0.27 0.60 0.06
(d) Phung et al. [45] 0.36 0.03 0.38 0.46 0.04
(e) StackedDiffusion [41] 0.43 0.02 0.42 0.57 0.07
(f) ShowHowTo 0.52 0.34 0.42 0.72 0.12
(g) Random 0.19 0.00 0.01 0.26 0.00
(h) Stable Diffusion [48$\dagger$ 0.70 0.03 0.44 0.84 0.03
(i) Copy of the input image 0.19 0.62 0.39 0.26 0.26
(j) Source sequences 0.50 1.00 0.56 0.60 1.00
$\dagger$ Generation not conditioned on the input image.
Table 1: Comparison with state-of-the-art on the ShowHowTo and the WikiHow datasets. Out of all the visual instruction generation methods, our method best follows the input prompts while being consistent with the input image.
Compared methods. We compare ShowHowTo to state-of-the-art methods for visual instruction generation as well as various baselines. For image-to-image methods (a-c), we generate the visual instructions sequence by iteratively using the last generated image as the input for the next step generation to achieve temporal consistency. (a) InstructPix2Pix [12] is trained to manipulate input images according to a text prompt by training on synthetic paired image data. In contrast, (b) AURORA [35] is trained on a manually curated dataset of image pairs from videos, while (c) GenHowTo [53] extracts the image pairs for training from instructional videos automatically.
Methods that generate image sequences (d-e) do not accept an input image, therefore, we apply the common input masking approach, where the first denoised frame of the sequence is replaced by the noised ground truth frame in each step of the generation. The (d) Phung et al. [45] method generates consistent sequences of visual instructions by attending to all frames in the sequence in the spatial attention layers. As it is a training-free method, we reimplement it and use it with the Stable Diffusion backbone [48]. (e) StackedDiffusion [41] is trained on WikiHow illustrated image sequences. It generates the image sequence as a single tiled image. Similarly to the related work, we evaluate all methods in zero-shot setup without finetuning.
Lastly, we show (g) Random lower bound and various naive baselines (h-i). (h) Stable Diffusion [48] is a text-conditioned generative method with no input image conditioning that generates each image independently, the (i) Copy baseline uses the input image as the output for any prompt. As an upper limit, (j) Source sequences uses the original dataset frames corresponding to the text prompts.
Quantitative results. We show the results of different methods on our ShowHowTo test set as well as on a subset of WikiHow sequences [62] in Table 1. The methods trained to perform localized edits (a-b) generate outputs fairly consistent with the input image (see the Scene Consistency metric), yet they fail to properly capture the instructions described by the text prompts (evidenced by the Step Faithfulness metric). On the other hand, methods for generating sequences of visual instructions (d-e) model the instructions well according to the input text prompts, but they perform poorly in scene consistency. In contrast, our method (f) generates visual instructions that are consistent with the input scene and correctly capture the action specified by the prompt. Our method even generates images that are more faithful to the input textual instructions than the dataset sequences (j) (see the Step Faithfulness metric). There are two reasons for this: objects in real images can appear small or occluded, impacting CLIP matching, and sometimes steps do not appear visually in the video.
Additionally, in Appendix E, we test our method in a zero-shot setup on the GenHowTo benchmark [53] and report additional quantitative metrics on the ShowHowTo dataset.
Figure 4: User study results. Win rates of the ShowHowTo method against baselines from pairwise forced decision user evaluations, divided into step, scene, and task. Values larger than 50% indicate ShowHowTo is preferred over the other methods (right).
User study. We present a user study with 9 participants evaluating sequences from 100 randomly sampled tasks from our test set. Each participant compared 50 ShowHowTo sequences with baselines using three criteria: (1) Step Faithfulness (Which sequence better follows the steps?), (2) Scene Consistency (Which sequence is more likely to come from the same video?), and (3) Task Faithfulness (Which sequence accurately depicts the instructions for the task of [task]?). As shown in Figure 4, ShowHowTo outperforms all baselines. Notably, users preferred our generations over sequences from source videos in 42% of cases for both step and scene metrics, which may be attributed to instructional videos not showing good views of steps at times and the high quality of our generation. Lower task faithfulness scores against the source sequences suggest room for improvement in future methods. More details are in Appendix D.
5.3 Ablations
We evaluate the key design decisions of our proposed method, i.e., the model conditioning, training data, and variable sequence length training, in the next paragraphs. Furthermore, additional performance analysis of the trained model is available Appendix E.
Text conditioning type Step Scene Task Average
Faithf. Consist. Faithf.
1 prompt (concatenated step prompts) 0.21 0.29 0.38 0.29
1 prompt (summarized step prompts) 0.20 0.30 0.40 0.30
1 prompt per step ($\tau_{0}=$ ‘an image’) 0.51 0.30 0.42 0.41
1 prompt per step (ShowHowTo) 0.52 0.34 0.42 0.43
Table 2: Ablation of step conditioning. The per-frame conditioning of ShowHowTo is instrumental in generating visual instructions faithful to the textual instructions.
Text model conditioning. We evaluate how different types of text conditioning affect model performance. For video models, it is common to provide a single text prompt for conditioning. However, visual instructions vary substantially from one another, possibly requiring different conditioning. We construct a single prompt for each sequence by concatenating all step prompts and by summarizing the step prompts using an LLM [19]. In Table 2, we show that our choice of separate prompt per step significantly outperforms both of the single prompt variants. Additionally, we demonstrate that using the free-form step description for $\tau_{0}$ outperforms the fixed prompt ‘an image’.
Model training data Step Scene Task Average
Faithf. Consist. Faithf.
WikiHow-VGSI [62] 0.55 0.12 0.30 0.32
HowToStep [37] 0.39 0.33 0.29 0.34
ShowHowTo (food videos only) 0.51 0.32 0.37 0.40
ShowHowTo 0.52 0.34 0.42 0.43
Table 3: Ablation of the training data as measured on the ShowHowTo test set. Our training dataset yields significant improvement over the manually curated WikiHow as well as the closely related HowToStep due to the quality of our instructions.
Training data. In Table 3, we analyze the impact of different training datasets by comparing our dataset with two related instructional datasets of similar scale and task coverage: (i) HowToStep [37], which contains automatically extracted video-text sequences from cooking videos, and (ii) WikiHow-VGSI [62], which consists of manually created image-text sequences from WikiHow articles, where the images primarily consist of digitally drawn illustrations. To train on the HowToStep dataset, we select the middle frame of each video segment as the visual instruction frame. We observe significantly worse performance caused both by the lack of precise instruction frame information as well as very noisy video segments (e.g., the dataset contains ‘Thank you for watching!’ segments which are not instructional). The performance is also worse when compared to the model trained only on the food-related ShowHowTo sequences that are extracted from the very same videos as the HowToStep sequences, indicating a superiority of our sequence extraction process. Training on WikiHow-VGSI yields higher Step Faithfulness score, likely due to the quality of manually matched images and prompts. However, the overall performance remains significantly below our approach, as the model primarily learns from illustrated images, resulting in less consistent scene generation.
Training sequence length Step Scene Task Average
Faithf. Consist. Faithf.
$\leq$ 4 steps 0.47 0.39 0.40 0.42
$\leq$ 8 steps (ShowHowTo) 0.52 0.34 0.42 0.43
$\leq$ 8 steps, randomly sampled 0.51 0.32 0.41 0.41
$=$ 8 steps 0.56 0.26 0.42 0.41
$\leq$ 16 steps 0.57 0.26 0.42 0.41
Table 4: Ablation of the training sequence length as measured on the ShowHowTo test set. We compare the performance of our model when trained on different sequence lengths.
Variable sequence length. Instructions are often of variable length, therefore, one of the key model requirements is to support generating image sequences of different lengths. While attention-based architectures allow for any sequence length, the question is how to train such a model. We test training the model on variable sequences of up to 4 frames, 8 frames, and 16 frames. We also train the model on sequences of length 8 only. In Table 4, we show that training on shorter sequences up to 4 frames results in high Scene Consistency but low Step Correctness, while training on sequences up to 16 frames is the opposite. This can be attributed to the fact that short sequences often keep the same background across the whole sequence, encouraging the model to preserve the background at the expense of the prompt. Longer sequences, on the other hand, have more variation of the background, e.g., as the task moves from the counter to the hob. The model is thus less likely to enforce the background during inference. Lastly, we also show that it is important to train always on consecutive sequences of visual instructions. If a subset of visual instructions from a video is sampled randomly (with the temporal ordering preserved), the scene consistency is decreased (Table 4, row 3).
Figure 5: Qualitative comparison using the input image (left) and the textual instructions (top) for the task of making a calzone. The images from the source video are shown in the first row. Except for ShowHowTo, methods either struggle to preserve the input scene or to produce coherent steps.
5.4 Qualitative Results
Figure 6: Qualitative results of our method for sequences from the test set. Given the input image (left) and the textual instructions (top), ShowHowTo generates step-by-step visual instructions while maintaining objects from the input image (e.g., the cooking pot and the ceramic bowl in rows one and four) as well as among generated images (e.g., glass bowl in the second row).
Qualitative results in Figure 1, Figure 6, and Figures 13, 14, and 15 in the appendix demonstrate the key strengths of the ShowHowTo model. It consistently preserves the scene as well as various objects, tools, and ingredients from the user-provided input image (e.g., pot and vegetables in Figure 6, first row). It dynamically adjusts the viewpoint to emphasize key actions. Similarly to vanilla text-to-image and text-to-video models, our model can also introduce plausible task-relevant objects (e.g., knives or bowls) if these objects are not present in the user-provided input image. Notably, the model effectively adapts human poses to demonstrate various object manipulations (e.g., flower arranging in Figure 6, third row).
Figure 5 shows qualitative comparisons with related methods. Methods for generating instructional sequences (StackedDiffusion [41] and Phung et al. [45]) fail to preserve the scene from the input image. For example, they generate blue, black, or wooden kitchen countertop instead of the glossy white one from the input image. On the other hand, image-to-image approaches (GenHowTo [53], AURORA [35], and InstructPix2Pix [12]) perform minimal edits and propagate errors through the output image sequence due to iterative generation (e.g., the persistent floating dough generated by the GenHowTo method).
Limitations. While our method can generate complex step-by-step visual instructions conditioned on the input image, it inherits the limitations of the models it is based on and introduces new limitations stemming from the novel source of training data. ShowHowTo model can struggle to maintain object states across many frames; for example, it can generate an image with raw meat after it was cooked in previous steps. Though the model often correctly generates common objects from instructional videos, for rare objects, such as electrical components, the model may generate objects in physically impossible configurations. Please see Figure 16 in the appendix for failure case examples.
6 Conclusion
This work explores, for the first time, generating environment-specific visual instructions to accomplish a user-defined task. We introduce a fully automated and scalable pipeline to create a dataset of 578K instructional image-text sequences from online videos, without requiring any manual supervision. Using this data, we train the ShowHowTo model to generate contextualized step-by-step visual instructions. Experiments demonstrate superior ability to generate accurate, scene-consistent instructional steps across various HowTo tasks, outperforming existing methods. We believe this work opens new avenues for personalized guidance in assistive technologies and step-by-step goal generation for robot planning.
Acknowledgements
We acknowledge VSB – Technical University of Ostrava, IT4Innovations National Supercomputing Center, Czech Republic, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (grant ID: 90254).
Research at the University of Bristol is supported by EPSRC UMPIRE (EP/T004991/1) and EPSRC PG Visual AI (EP/T028572/1). Prajwal Gatti is partially funded by an uncharitable donation from Adobe Research to the University of Bristol.
This research was co-funded by the European Union (ERC FRONTIER, No. 101097822 and ELIAS No. 101120237) and received the support of the EXA4MIND project, funded by the European Union’s Horizon Europe Research and Innovation Programme, under Grant Agreement N° 101092944. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.
Appendix
In the appendix, we first provide dataset collection details and show examples from our dataset in Section A. Then, in Section B, we compare our dataset to other related works and discuss their differences. We provide implementation details in Section C and evaluation details and analysis of our metrics in Section D. In Section E, additional quantitative results are provided, and in Section F, we show a large variety of qualitative results.
Appendix A Dataset Collection Details
Speech transcription: YouTube ASR vs. WhisperX. The original video transcripts released with the HowTo100M dataset, generated through YouTube ASR, are known to contain errors [25, 37]. Following Li et al. [37], we employ WhisperX [4] to obtain higher-quality transcripts for narrated instructional videos. As shown in Figure 18, WhisperX provides improved transcription quality over YouTube ASR through better punctuation, fewer transcription errors (e.g., “Put some aluminum foil in there” vs. YouTube ASR’s incorrect “it’s a moment of oil in there”), and more accurate sentence segmentation with timestamps. These improvements are essential for our subsequent dataset creation steps, which rely on accurate transcription.
Details on filtering of irrelevant videos. To address the presence of non-instructional content in HowTo100M, we prompt Llama 3 [19] (Llama-3.1-8B-Instruct) to identify whether videos are instructional or not based on their transcript excerpts. The full prompt used is shown in Figure 19.
We validate our approach for video filtering by manually annotating a balanced validation set of 200 videos. Our approach achieves the false positive rate of 5% and the false negative rate of 12%, i.e., 95% of non-instructional videos spuriously present in the HowTo100M dataset are filtered out by our automatic approach. Applied to the full HowTo100M dataset, we identified 847K (68.4%) instructional videos and filtered out 391K (31.6%) non-instructional content, including product reviews (e.g., “Houseplant Unboxing | Steve’s Leaves”), entertainment videos (e.g., “Don McLean - American Pie (with Lyrics)”), and personal vlogs (e.g., “Driving to the West, an RV lifestyle vlog”).
Figure 7: Frame matching comparison across CLIP, SigLIP, and DFN-CLIP (used in our work). For each method, the figure shows the best matching frame to the instructional text (shown below).
Figure 8: Samples from the ShowHowTo dataset. Each sample (row) is a sequence of textual instructions (top) and the associated visual instruction images (bottom).
18.05 - 19.53 Put aluminum foil in the pressure cooker.
28.86 - 29.70 Create two little rings using the aluminum foil.
38.50 - 40.42 Place the potatoes in the pressure cooker, above the water.
44.99 - 50.23 Add 1.5 cups of water to the pressure cooker.
59.85 - 62.05 Cook the potatoes for 15 minutes, depending on their size.
76.08 - 80.48 ? Remove the potatoes from the pressure cooker and serve.
4.00 - 12.00 ? Put potatoes in the pressure cooker and create steaming rings.
17.00 - 25.00 ? Add one and a half cups of water to the pressure cooker.
19.00 - 27.00 Place aluminum foil in the pressure cooker.
78.00 - 86.00 It doesn’t heat up your kitchen.
83.00 - 91.00 It’s really good for in the summer when you want to have
something like a baked potato.
86.00 - 94.00 Up anything else.
91.00 - 99.00 When they’re done, you just pop
96.00 - 104.00 It’s just going to be steamed.
96.00 - 104.00 So there it is.
an instruction with correct approximate timestamp ?an instruction with incorrect timestamp not an instruction
Figure 9: Comparison between textual instructions extracted by our method (left) and the textual instructions from HowToStep [37] (right) for the same randomly chosen ‘How to bake a potato in the pressure cooker’ video. The original transcript used to produce our instructions is shown in Figure 18. Our method correctly identifies the key steps in the narrations and summarizes them in step-by-step instructions. On the other hand, the HowToStep data often contain steps that are not instructions.
Step extraction details. Our approach uses Llama 3 to extract the steps from video narrations. This contrasts the related work [53], which extracts steps from video frames and uses an image-captioning model to generate step captions. We found that the latter approach resulted in captions that were too brief, high-level, and lacked sufficient detail to differentiate adjacent steps.
Figure 20 illustrates our Llama 3 prompt for extracting instructional steps from video narrations, including the few-shot examples used (truncated to fit into a page; complete prompts are available on the project website). We processed videos under 10 minutes long, as longer transcripts exceeded the context limit of Llama 3 and also degraded output quality. The model is instructed to generate temporally ordered steps with approximate timestamps using video transcripts. We filter out malformed results with non-temporally ordered steps or incomplete descriptions. Examples of extracted steps are shown in Figure 8.
Cross-modal frame alignment details. For each extracted step, we find its matching frame in the video using DFN-CLIP [21] (DFN5B-CLIP-ViT-H-14-378), restricting the search to frames within the step’s time bounds generated by Llama 3. We formulate this as a dynamic programming problem to find optimal frame-text pairs while preserving temporal order and maximizing alignment scores.
Although the temporal boundaries generated by Llama 3 are rarely incorrect, narrations do not always align with respect to video frames and can occur slightly before/after the visuals. To account for the well-known temporal misalignment [25, 37], we expand the temporal boundaries by a fixed duration of $\epsilon$ seconds, increasing the search space for frame alignment. Through analysis on a small validation set of manually annotated videos with precise step boundaries, we found $\epsilon=15$ seconds yielded a good balance of precision (the correct frame was selected by DFN-CLIP) and recall (the correct frame was inside the searched interval).
We use DFN-CLIP over related contrastive models such as CLIP [47] and SigLIP [66] as we found it superior in certain cases for matching video frames to instructional steps. We observed CLIP and SigLIP often exhibited limitations such as incorrect object state identification (e.g., unroasted peppers, uncut prosciutto in Figure 7, rows 1-2) and tendency to select blurry or transition frames (Figure 7, rows 2-3). We quantify the quality of the automatically selected keyframes by a small user study. We manually annotated keyframes for 100 steps and asked humans to blindly select whether the manually or automatically selected frame better corresponds to the text instruction. DFN-CLIP was preferred in 18% of cases, human annotation was preferred in 36% of cases, remaining 46% of cases were a tie. This indicates fairly decent alignment with human judgment.
Dataset Statistics. We analyze the ShowHowTo dataset statistics in Figures 10 and 11. The dataset encompasses 25K tasks across diverse categories, including Food and Entertainment, Hobbies and Crafts, etc., derived from HowTo100M’s task hierarchy. Figure 10 (left) shows the distribution of these categories in our dataset. Each sample contains an average of 7.7 steps, with 11.37 words per step (Figure 10, right). The word clouds in Figure 11 showcases common verbs of physical actions like remove, add, and make, alongside various household objects and materials used in everyday tasks.
Figure 10: Statistics of the ShowHowTo dataset. Left: Distribution of task categories in ShowHowTo dataset. Top-right: Distribution of the number of steps per sequence. Bottom-right: Distribution of the number of words per sentence.
Figure 11: Word cloud visualizations of verbs (left) and nouns (right) in the textual instructions of ShowHowTo dataset.
Dataset Source Manually Task Visual Scale # Tasks Avg. Visual Type Annotation Type
Curated Domain Domain Steps / Seq.
CrossTask [70] YouTube Open Open 4.7K 18 7.4 Video Segments Categorical
COIN [54] YouTube Open Open 10K 180 3.9 Video Segments Categorical
Ego4D Goal-Step [51] Ego4D Open Egocentric 717$\ddagger$ 80 23.3 Video Segments Instructions
LEGO [36] Ego4D & EPIC-K. Open Egocentric 147K - 1.0 Key Frames Instructions
AURORA [35] Multiple Open Open 289K - 1.0 Key Frames Instructions
GenHowTo [53] COIN & ChangeIt Open Open 45K 224 2.0 Key Frames Captions
HowToCaption [50] HowTo100M Open Open 1.1M 23.6K 18.5 Video Segments Captions
HT-Step [2] HowTo100M Cooking Open 18K 433 5.9 Video Segments Instructions
HowToStep [37] HowTo100M Cooking Open 312K 14.2K 10.6 Video Segments Instructions
[…]
ShowHowTo HowTo100M Open Open 578K 25K 7.7 Key Frames Instructions
$\dagger$ Some examples in the dataset are real photos.  $\ddagger$ The subset with step annotations.
Table 5: Comparison of instructional datasets. Scale refers to the number of instruction sequences of image-text pairs. Annotation Type describes the nature of text captions. Visual Type indicates the format of visual content.
Appendix B Relation to Existing Datasets
We show the comparison of related instructional datasets in Table 5. Early datasets like CrossTask [70] and COIN [54] are manually curated but small in scale with categorical instruction annotations. While recent datasets like HowToCaption [50] have expanded significantly (1.1M sequences), they provide generic captions rather than instructional text annotations. Specialized datasets exist for egocentric domains (LEGO [36], Ego4D Goal-Step [51]) and single-step instructions (AURORA [35], GenHowTo [53]), but their narrow scope limits general applicability.
The two most comparable large-scale multi-step instructional datasets to our dataset are WikiHow-VGSI [62] (100K sequences) and HowToStep [37] (312K sequences). WikiHow-VGSI, composed of image-step pairs extracted from WikiHow articles, predominantly contains digitally drawn illustrations rather than real photos, making it unsuitable for realistic image generation and very difficult to scale. On the other hand, HowToStep [37], similarly to us, leverages HowTo100M videos, but its focus is solely on the cooking domain, and it contains only temporal video segments instead of individual representative images. Lastly, the provided segments are of variable quality, often not being instructional (see Figure 9 for an example). In comparison, the ShowHowTo dataset is both larger in scale and contains more diverse tasks with higher-quality steps. In Section 5.3, we train our model on both WikiHow-VGSI and HowToStep, showing that training on our dataset results in models with substantially better generation capabilities.
Appendix C Implementation Details
ShowHowTo implementation details. We initialize our model with pretrained weights of DynamiCrafter for image animation [56] and train the model for approximately 100K steps on four AMD MI250x multi-chip modules (8 GPUs) with a total of 512 GB of VRAM (64 GB per GPU). We use the batch size of 16 image sequences at the resolution of 256$\times$256 pixels. The image sequence length is variable, ranging between two and eight. We train the model using AdamW optimizer with a learning rate of $2\cdot 10^{-5}$. The training takes approximately 48 hours. During inference, we generate visual instruction sequences using the DDIM sampler with 50 denoising steps. Both the inference and training code, along with trained model weights are publicly available at https://soczech.github.io/showhowto/.
Implementation details of the related methods. In our evaluation of the related methods, we use the official implementations of GenHowTo, AURORA, and InstructPix2Pix provided by the respective authors. In the case of GenHowTo, we use the action model weights. For StackedDiffusion [41], we use the official implementation modified to replace the first generated image with the noised version of the input image in each denoising step, to enable input image conditioning. The work of Phung et al. [45] does not provide the official implementation at the time of writing; therefore, we reimplement it using the Stable Diffusion 2.1 model. For the step similarity matrix, we use the matrix with all values set to one, and similarly to the StackedDiffusion model, we replace the first generated image with the noised version of the input image in each denoising step.
Figure 12: Correlation of our metrics with human preference. User study results (Win Rate %) highly correlate with our three metrics—Step Faithfulness, Scene Consistency, and Task Faithfulness.
Appendix D Evaluation Details
As described in Section 5.1, we evaluate our method using three metrics: Step Faithfulness, Scene Consistency, and Task Faithfulness. DFN-CLIP [21] (DFN5B-CLIP-ViT-H-14-378) is used for the computation of the Step Faithfulness and Task Faithfulness metrics. Scene Consistency is computed with the averaged spatial patch features of DINOv2 [44] (dinov2_vitb14_reg). All metrics are first averaged per sequence before being averaged across the test set to account for the variable sequence length.
We verify how well the used metrics correlate with human preference as measured by our user study (see Section 5.2). For each evaluation metric, we plot the performance of all other models against the win rate compared to the ShowHowTo model from the user study. The results can be seen in Figure 12, we also show the line of best fit and $R^{2}$ value. We find a high correlation across all metrics, especially so for Scene Consistency and Task Faithfulness, confirming that our metrics serve as a good proxy for human preference.
Figure 13: Qualitative results of our method for sequences from the test set. Our model can generate both short instructional sequences (as shown here) as well as long or very long sequences shown in Figures 14 and  15.
Appendix E Additional Quantitative Results
[…]
InstructPix2Pix [12] 37.8
AURORA [35] 23.2
GenHowTo [53] 28.3
Phung et al. [45] 27.8
StackedDiffusion [41] 34.6
ShowHowTo 12.4
Table 6: Comparison with state-of-the-art using the FID score on the ShowHowTo test set. ShowHowTo model significantly outperforms all related methods.
Method Accac
Stable Diffusion [48] 0.51
Edit Friendly DDPM [31] 0.60
InstructPix2Pix [12] 0.55
GenHowTo [53] 0.66
ShowHowTo 0.72
Table 7: Zero-shot evaluation on the GenHowTo dataset according to the GenHowTo protocol [53]. The ShowHowTo model outperforms the prior state-of-the-art with no fine-tuning on the GenHowTo train set.
FID results. We also evaluate all methods using the FID score in Table 6. For each input image $I_{0}$ and the textual instructions $\{\tau_{i}\}_{i=0}^{n}$ from the ShowHowTo test set, we generate the sequence $\{\hat{I}\}_{i=1}^{n}$ of visual instructions. For each method, its FID score is computed between its generated sequences and the source visual instruction sequences $\{{I}_{i}\}_{i=1}^{n}$ from the test set. Our method generates images that better match the distribution of real visual instruction sequences.
Zero-shot evaluation on the GenHowTo dataset. In GenHowTo [53], the authors propose to evaluate generative methods in a downstream application, where the method generates images of various classes that are then used to train a simple classifier. The performance of this classifier is then computed on the real set of images. We show the results of our method in Table 7. We evaluate our model as is, i.e., trained on the ShowHowTo dataset without any additional training or fine-tuning. We report the action accuracy metric $Acc_{ac}$, which evaluates whether our generated visual instruction images can be used for downstream application of classifying actions. Note that this metric is image-based and does not evaluate sequences. Nonetheless, ShowHowTo improves over the previous state-of-the-art from [53] by 6 percentage points.
Number of generated frames
1 2 3 4 5 6 7
Step Faith. 1.00 0.72 0.59 0.52 0.50 0.50 0.51
Scene Consist. 0.20 0.29 0.43 0.42 0.34 0.36 0.31
Task Faith. 0.30 0.49 0.48 0.45 0.42 0.46 0.37
Table 8: Model’s performance for different lengths of the generated sequence. The performance is fairly similar across different lengths, with a decrease observed for longer sequences.
Variable sequence length generation analysis. Our model generates sequences of variable length. We analyze the performance of the model for various sequence lengths in Table 8. Except for the degenerative case with one frame, the performance is fairly similar across different lengths, with a decrease observed for longer sequences.
Task category Step Scene Task Average
Faithf. Consist. Faithf.
Cars & Other Vehicles 0.34 0.37 0.58 0.43
Education and Communications 0.37 0.55 0.24 0.39
Food and Entertaining 0.67 0.24 0.40 0.44
Health 0.29 0.56 0.46 0.44
Hobbies and Crafts 0.39 0.46 0.41 0.42
Holidays and Traditions 0.56 0.35 0.52 0.48
Home and Garden 0.40 0.41 0.38 0.40
Pets and Animals 0.48 0.41 0.39 0.43
Sports and Fitness 0.36 0.42 0.55 0.44
Table 9: Model’s performance for different task categories. Task categories with only a few sequences not shown.
Per-task performance analysis. We report the performance of our model in various HowTo100M task categories in Table 9. We observe that the performance is significantly dependent on the task distribution in the dataset. In detail, for the Step Faithfulness metric, the best performance is achieved in cooking tasks because of plentiful training data and clear, visually distinct steps. On the other hand, for Scene Consistency, the cooking tasks perform the worst as the matching is done across the whole dataset, where a large portion is cooking with many similar scenes and frames. Additionally, the cooking tasks contain many close-up scenes without any scene background that can be used for matching the generated images to the correct sequence. The best Scene Consistency is achieved for tasks in the Health and the Education categories due to the uniqueness of the sequences in those categories.
Figure 14: Additional qualitative results of our method for sequences from the test set. Given the input image (left) and the textual instructions (top), ShowHowTo generates step-by-step visual instructions while maintaining objects from the input image.
Figure 15: Qualitative results of our method for sequences from the test set. Our model can generate long sequences of visual instructions while being consistent with the input image and the text prompts.
Figure 16: Failure modes. The model can struggle with rare objects and tools (left), or it can fail to update object states after state-changing actions (right).
Appendix F Additional Qualitative Results
Additional qualitative results. We show additional qualitative results in Figures 13, 14, and 15. We show our method can correctly generate sequences of visual instructions according to the input images and prompts. In Figure 15, we demonstrate our method can generate long instructional sequences while preserving consistency with the input image. Additionally, in Figure 13, we show our model can also generate shorter sequences.
Additional qualitative comparison with related work. We show additional comparison with related work on the task of creating paper flowers in Figure 17. We can see our method not only correctly captures the scene, which is not the case for the method of Phung et al. [45] and StackedDiffusion [41], but the model faithfully follows the input prompts, generating useful visual instructions for the user.
Failure modes. We show some limitations of our method, as described in the main paper, in Figure 16. Our model can struggle with objects that are not common in the training data, such as engine cylinders and tools such as razor blades. Additionally, the model can make errors in scenarios where object states need to be tracked and updated across multiple frames, such as after cooking meat, the meat’s state must be changed from raw to cooked, etc.
Figure 17: Additional qualitative comparison using the input image (left) and the textual instructions (top) for the task of making a cardboard flower with glitter. Only ShowHowTo can produce convincing steps while preserving the input scene.
(keyval) Package keyval Error: breaklines undefinedSee the keyval package documentation for explanation.I can’t help
  0.50 -   2.60: Hi, it’s Matthew in his pressure cooker again.
  2.60 -   6.09: And today I’m going to make baked potatoes in the
                 pressure cooker.
  6.09 -  11.71: Obviously they’re not baked potatoes, just an easy way
                 to make something like a baked potato.
 11.71 -  13.29: It’s basically going to be steaming them.
 13.29 -  18.05: So I use my aluminum foil trick.
 18.05 -  19.53: Put some aluminum foil in there.
 19.53 -  27.84: Now a lot of pressure cookers do come with standoffs so
                 that you can put things in where they’re sitting above
                 the water.
 28.86 -  29.70: I’m going to make two little rings.
 29.70 -  38.50: I’m going to take my potatoes, put them in so that
                 they’re going to be steamed nicely.
 38.50 -  40.42: Because you don’t want to actually boil them.
 40.42 -  42.28: You want to steam them.
 42.28 -  44.99: And I’ve got my cup and a half of water.
 44.99 -  50.23: I’ll just dump that in.
 50.23 -  52.27: See what I’ve got in here?
 52.27 -  53.47: So they’re just sitting there.
 53.47 -  54.31: They’re above the water.
 54.31 -  57.49: Now, the amount of time varies quite a bit.
 59.85 -  62.05: For these ones I’m going to put them in for about 15
                 minutes.
 62.05 -  63.76: Smaller potatoes maybe a little bit less.
 63.76 -  65.02: Larger potatoes a little bit more.
 65.02 -  68.52: The nice thing about this is it doesn’t tie up anything
                 else.
 68.52 -  69.90: It doesn’t heat up your kitchen.
 69.90 -  73.88: It’s really good for in the summer when you want to have
                 something like a baked potato.
 73.88 -  75.48: It’s just going to be steamed.
 75.48 -  76.08: So there it is.
 76.08 -  76.94: 15 minutes.
 76.94 -  80.48: When they’re done you just pop them out and eat them
                 like a normal baked potato.
 80.48 -  81.47: I hope you find this useful.
 81.47 -  87.93: If you want to hear more ideas or have any questions
                 leave a comment, send me an email and I’ll see what I
                 can do for you.
 89.45 -  92.67: I hope you’re enjoying your pressure cooker as much as
                 I’m enjoying mine.
 92.67 -  93.92: Bye.
 93.92 -  96.84: And here we go.
 96.84 - 101.70: I’ll pull it out.
101.70 - 104.60: There is the baked potato or steamed potato actually.
104.60 - 107.52: It’s nice and it’s pretty dry on the outside.
108.55 - 110.37: Just cut it open.
110.37 - 112.43: It’s nice and soft.
112.43 - 115.07: Well cooked on the inside.
115.07 - 117.41: Great to cook it up however you’re going to make a meal.
117.41 - 119.35: If you’re going to eat it like a traditional baked
                 potato.
119.35 - 121.63: If you’re going to use it for potato salad or whatever.
121.63 - 123.79: I hope you enjoyed it.
123.79 - 126.75: If you want to see any other ideas, check my channel.
126.75 - 128.35: See what other things I’ve got posted.
128.35 - 134.16: If you’ve got ideas that you don’t know how to do,
                 send me an email or leave a comment and I’ll see what I
                 can do.
134.16 - 134.70: Hope you enjoyed it.
134.70 - 135.04: Bye.
(keyval) Package keyval Error: breaklines undefinedSee the keyval package documentation for explanation.I can’t help
hi it’s math units pressure cooker again
and today I’m going to make baked
potatoes in the pressure cooker
obviously they’re not baked potatoes
just an easy way to make something like
a baked potato it’s basically going to
be steaming them so I use my aluminum
foil trick yeah it’s a moment of oil in
there now a lot of pressure cookers do
come with standoffs but you can cook
eggs in where they’re sitting above the
water and make two little rings I’m
going to take my potatoes put them in so
they’re going to be steamed nicely
because you don’t want to actually boil
them on it you want to see them and go a
cup and half water I’ll just dump that
in see see what I’ve got in here so
they’re just sitting there they’re above
the water now the amount of time varies
quite a bit for these ones I’m going to
put them in for about 50 minutes smaller
potatoes may be a little bit less larger
potatoes a little bit more a nice thing
about this is it doesn’t tie up anything
else doesn’t heat up your kitchen it’s
really good for in the summer when you
want to have a something like a baked
potato it’s just going to be steamed so
there it is 15 minutes when they’re done
you just pop them out and eat them like
a normal baked potato I hope you find
this useful if you want to hear more
ideas or have any questions leave a
comment send me an email and I’ll I’ll
see what I can do here enjoy your I hope
you’re enjoying your pressure cooker as
much as I’ve enjoyed mine bye and here
we go hold it there is Becky or steamed
potato actually placing some pretty dry
on the outside just cut it open it’s
nice and nice this off well cooked on
the inside great to cook it up however
you’re going to make me know if you need
it like a traditional baked potato we’re
going to use it for potato salad or
whatever I hope you enjoyed it if you
want to see any other ideas so, check my
channel see the other things I’ve got
posted you’ve got ideas so that you
don’t know how to do send me an email or
leave a comment and I’ll see what
you do hope you enjoyed it I
Figure 18: Comparison between WhisperX [4] speech transcription (left) and YouTube ASR (right) for the same ‘How to bake a potato in the pressure cooker’ video. In contrast to YouTube ASR, WhisperX can correctly split the narrations into individual sentences. It also makes significantly fewer errors; for example, it correctly recognizes that the potatoes should be cooked for 15 minutes, not 50 minutes (timestamp 59.85).
Prompt for filtering non-instructional videos Based on the following video title and partial transcript segment, determine if the video is instructional in nature, where “instructional” means it involves actively demonstrating or teaching how to perform a specific task or activity with physical steps (e.g., cooking a recipe, repairing something, crafting, etc.). Respond with ‘Yes’ if the video is actively demonstrating or teaching how to perform a specific task, or ‘No’ if it is not. Then provide a single sentence explanation.
Examples of instructional videos:
• How to Bake a Chocolate Cake
• Repairing a Leaky Faucet
• Learn to Knit a Scarf
Examples of non-instructional videos:
• Discussing Fashion Trends
• Product Reviews and Opinions
• A Vlog of My Daily Life
Example 1:
Video Title: Red Dead Redemption 2 - Herbert Moon and Strange man Eastereggs In Armadillo [SPOILERS]
Video Transcript: “oh you’re back I feared the worst it’s all here waiting for you who’s that I don’t know it’s just a little portrait somebody gave me once I always quite liked it why no reason just seem familiar anyway this area is closed to the public if you want to shop here you better act right move you long streak of piss who do you think you are for God’s sake get out you degenerate you blew it get out of my store if you don’t leave there will be problems okay okay stay calm oh you’ll (…)”
Is this video actively demonstrating or teaching how to perform a specific task? No
Explanation: The video is not actively demonstrating or teaching how to perform a specific task; it appears to be showcasing or discussing Easter eggs in the game Red Dead Redemption 2.
Example 2:
Video Title: Fantastic VEGAN Cupcakes with Raspberry Frosting
Video Transcript: “hey there I’m chef Annie and tomorrow is Valentine’s Day so we are making some extra special cupcakes for this occasion can you believe that we have not made cupcakes on this channel it’s about time so today I’m going to show you how to present these cupcakes so they look impressive and absolutely beautiful so enough copy let’s cook it so we’re going to start by mixing together our wet ingredients (…)”
Is this video actively demonstrating or teaching how to perform a specific task? Yes
Explanation: The video actively demonstrates and teaches how to make vegan cupcakes with raspberry frosting, as indicated by the detailed steps and instructions given by the chef.
Example 3:
Video Title: How To: Piston Ring Install
Video Transcript: “hey it’s Matt from how to motorcycle repair comm just got done doing a top end on a YZF 250 or yz250 F and I thought I’d do a quick video on how to install a piston ring the easy way now I’ve done this in the past too but most people will take the ends here and spread it and put it on but you can potentially damage the ring so an easier way to do that is just to take this right here incident in the groove that you need then you bend one up (…)”
Is this video actively demonstrating or teaching how to perform a specific task? Yes
Explanation: The video is actively demonstrating or teaching how to install a piston ring, which is a specific task.
Example 4:
Video Title: Best gas weed eater reviews Husqvarna 128DJ with 28cc Cycle Gas Powered String Trimmer
Video Transcript: “guys i’m shanley today i’m going to tell you about this straight shaft gas-powered trimmer from husqvarna this trimmer runs on a 28 CC two cycle engine it features 1.1 horsepower and a three-piece crankshaft it also has a smart start system as well as an auto return to stop switch and this trimmer is air purge design for easier starting it has a 17 inch cutting path (…)”
Is this video actively demonstrating or teaching how to perform a specific task? No
Explanation: This video is reviewing the features of a gas-powered trimmer rather than actively demonstrating or teaching how to use it.
Now, determine if the following video is instructional in nature: Video Title: {Input video title} Video Transcript: {Input video transcript} Is this video actively demonstrating or teaching how to perform a specific task?
Figure 19: Prompt used for filtering non-instructional videos using Llama 3. Transcript excerpts are truncated for clarity, the full prompt is available on the project website.
Prompt for step extraction from an instructional video Below are transcripts from YouTube instructional videos and their corresponding extracted steps in a clear, third-person, step-by-step format like WikiHow. Each step is concise, actionable, and temporally ordered as they occur in the video. The steps include start and end timestamps indicating when the steps are carried out in the video. Follow this format to extract and summarize the key steps from the provided transcript.
Example 1:
YouTube Video Title: “BÁNH TÁO MINI - How To Make Apple Turnovers | Episode 11 | Taste From Home
YouTube Video Transcript:
00.87 - 07.79: “Hey little muffins, today we will make together a super easy, quick and delicious apple turnovers."
07.79 - 09.35: “40 minutes for all the process."
09.35 - 11.95: “Seriously, can someone deny them?"
11.95 - 13.63: “Ok, let’s begin."
13.63 - 18.82: “First of all, combine the apple cubes, lemon juice, cinnamon and sugar in a bowl."
26.69 - 29.59: “Mixing, mixing, mixing."
29.59 - 32.62: “Apple and cinnamon always go perfectly together."
32.62 - 43.52: “Now using a round cutter or glass like me, cut 15 rounds from the pastry sheet."
57.86 - 64.99: “Here comes the fun part."
64.99 - 69.97: “Spoon about 2 teaspoons apple mixture in the center of one round."
69.97 - 74.41: “Using your fingers, gently fold the pastry over to enclose filling."
88.47 - 104.48: “After that, use a fork and press around the edges to seal and make your apple turnovers look more beautiful."
104.48 - 105.84: “This is how it looks like."
109.99 - 113.53: “I will show you one more time to make sure that you understand the technique."
113.53 - 117.20: “And if you still find my apple turnovers too ugly, I’m really sorry."
Extracted Steps: [{ "WikiHow Title": "How to Make Apple Turnovers" }, { "steps": [ { "step": 1, "instruction": "Combine apple cubes, lemon juice, cinnamon, and sugar in a bowl.", "start_timestamp": 13.63, "end_timestamp": 18.82 }, { "step": 2, "instruction": "Mix the ingredients thoroughly.", "start_timestamp": 26.69, "end_timestamp": 29.59 }, { "step": 3, "instruction": "Cut 15 rounds from the pastry sheet using a round cutter or a glass.", "start_timestamp": 32.62, "end_timestamp": 43.52 }, { "step": 4, "instruction": "Spoon about 2 teaspoons of the apple mixture into the center of one round.", "start_timestamp": 64.99, "end_timestamp": 69.97 }, { "step": 5, "instruction": "Gently fold the pastry over to enclose the filling using your fingers.", "start_timestamp": 69.97, "end_timestamp": 74.41 }, { "step": 6, "instruction": "Press around the edges with a fork to seal and beautify the turnovers.", "start_timestamp": 88.47, "end_timestamp": 104.48 }, { "step": 7, "instruction": "Repeat the technique until all turnovers are formed.", "start_timestamp": 109.99, "end_timestamp": 113.53 }, { "step": 8, "instruction": "Lightly beat one egg in a small bowl.", "start_timestamp": 151.62, "end_timestamp": 157.46 }, { "step": 9, "instruction": "Egg wash the apple turnovers to give them a gorgeous light brown color after baking.", "start_timestamp": 157.46, "end_timestamp": 164.10 }, { "step": 10, "instruction": "Bake the apple turnovers at 180°C for 18-20 minutes until golden.", "start_timestamp": 164.10, "end_timestamp": 174.87 }, { "step": 11, "instruction": "Enjoy the freshly baked apple turnovers.", "start_timestamp": 178.17, "end_timestamp": 185.55 }]}] Example 2: (…)
Now, extract the steps from the following transcript:
YouTube Video Title: {Input video title}
YouTube Video Transcript: {Input video transcript}
Extracted Steps:
Figure 20: Prompt used for generating instructional steps with start-end timestamps using Llama 3. The prompt is truncated for clarity, the full prompt is available on the project website.